### v2

# CaMML: Context-Aware Multimodal Learner for Large Models

Abstract
In this work, we introduce Context-Aware MultiModal Learner (CaMML), for tuning large multimodal models (LMMs). CaMML, a lightweight module, is crafted to seamlessly integrate multimodal contextual samples into large models, thereby empowering the model to derive knowledge from analogous, domain-specific, up-to-date information and make grounded inferences. Importantly, CaMML is highly scalable and can efficiently handle lengthy multimodal context examples owing to its hierarchical design. Based on CaMML, we have developed two multimodal models, CaMML-7B and CaMML-13B, that have shown exceptional performance across an array of benchmark datasets for multimodal tasks. Remarkably, CaMML-13B achieves the state-of-the-art performance on over ten widely recognized multimodal benchmark datasets, surpassing LLaVA-1.5 (13B) with a noticeable margin, without integration of any external resources. Moreover, we have conducted extensive ablative studies to inspect the inner workings of CaMML and performed qualitative analyses to showcase its effectiveness in handling real-world challenging cases.
[…]
1 Introduction
Figure 1: CaMML achieves the state-of-the-art performance on a number of multimodal benchmarks, outperforming LLaVA-1.5 and many other large multimodal models.
Recently, large multimodal models (LMMs) [32, 61, 54, 31, 66, 53, 3] have demonstrated remarkable performance in a variety of tasks, including but not limited to visual question answering, image captioning, visual grounding, visual-language reasoning, optical character recognition, and visual entailment. Notably, in certain benchmark assessments, these multimodal foundation models have even exceeded human-level performance [63, 32].
Despite the impressive performance, their ability to make inferences is constrained by the knowledge encoded in the model parameters. The inflexible design of these models makes it challenging for them to generalize from contextual examples. For instance, LLaVA-1.5 falls short in processing multiple images and attributes it to the lack of corresponding instruction-tuning training data [31]. Nevertheless, learning and making inferences through contextual examples are fundamental elements of our cognitive processes. Human beings frequently tackle intricate problems by relying on past experiences and identifying analogous situations. Taking inspiration from the cognitive process, we hypothesize that empowering large multimodal models with the capability to perceive and derive insights from analogous contextual examples can significantly streamline the inference process and lead to more precise predictions. Nonetheless, the means to replicate the human cognitive processes in LMMs remain unclear.
As such, our goal is to empower multimodal foundational models to harness context-aware learning, thereby enhancing their ability to comprehend and adapt to previously unseen examples. Identifying relevant context examples is relatively straightforward; this process can be facilitated using multimodal embedding models like ImageBind [15] or CLIP [43]. This approach simulates the act of recalling similar situations from past experiences. Yet, effectively and efficiently integrating the identified similar context samples into large models poses challenges, particularly given the potential variability in the number of context samples and interleaved modalities, resulting in lengthy and heterogeneous context input.
To this end, we propose a context-aware multimodal learner, dubbed as CaMML, for LMMs. CaMML acts as a crucial intermediary between the contextual examples and a large language model (LLM). Our approach is structured hierarchically, where the initial level establishes connections between the text and image modalities for each example through cross-attention mechanisms. This integration of text and image information enables a deeper understanding of the interleaved context. Following this, another module takes the outputs of the first level and performs cross-attention between the contextual information and a predefined set of learnable, fixed-length tokens. The resulting output from this level is then used as fixed-length input for the LLM, allowing the model to leverage the refined and context-aware information to perform complex multimodal understanding tasks. To summarize, we make the following contributions:
  • We propose CaMML, a context-aware multimodal learning approach for finetuning multimodal models. CaMML is lightweight and can be applied to process extremely long multimodal context samples;
  • With CaMML, we have developed two multimodal models, CaMML-7B and CaMML-13B. These models have achieved state-of-the-art performance across a diverse range of benchmarks encompassing various multimodal tasks, all without the need for external data integration;
  • We conduct comprehensive model analyses and case studies to examine the internal mechanisms of CaMML and showcase how the proposed model can effectively handle real-world challenging cases.
2 Related Work
Large Multimodal Models
The success of LLMs has sparked a burgeoning interest in scaling up multimodal models. One prevalent strategy involves the incorporation of vision encoders, such as ViT [13] and CLIP, into existing LLMs (e.g., LLaMA [52], Vicuna [65]). For example, LLaMA-adapter [61] inserts learnable adaption prompts into LLaMA and make it to follow multimodal instructions for multimodal reasoning. BLIP2 [28] bridges different modalities via Q-Former and employs a two-stage training method to bootstrap both representation learning and LLM (e.g., OPT [62], Flan-T5 [11]) generative learning capabilties. LLaVA [32, 31] projects encoded image features to text token space using linear layers and shows state-of-the-art performance on a variety of multimodal tasks. Another noteworthy series of strategy is to unify multimodal input data and pretrain the model from scratch, exemplified by OFA [54], Perceiver [19, 20], Uni-Perceiver [67], Unival [47], and Unified-IO [35]. These models map images, text, and other modalities into the same IO space and use a causal language model objective for model training.
Multimodal Few-shot Learning
Extensive research has been conducted on learning from a limited number of multimodal examples. Among them, Flamingo [3] introduces gated xatten-dense layers to establish cross-modal interactions between visual input and text input for few-shot learning. Frozen [53] trains a vision encoder to produce a sequence of image embeddings and input it to frozen language models for multimodal tasks. An alternative approach for tackling multimodal few-shot learning involves leveraging retrieval augmented generations (RAG) [26, 7, 44, 23]. This technique enables the models to access external knowledge repositories, databases, or structured data sources, allowing them to access the up-to-date as well as domain-specific expertise when crafting responses. Recently, there has been a surging interest for harnessing RAG within the realm of multimodal models [9, 55, 56, 33]. Amalgamating RAG with multimodal models helps provide contextually-rich responses and opens up exciting possibilities across domains such as image captioning [55], image generation [56], etc. Among them, Chen et al. [9] propose MuRAG, a model of size 527M, for multimodal-QA. MuRAG concatenates the visual embeddings (extracted by ViT) and text word embeddings of retrieved multimodal image-text pairs and requires the language model to handle a very lengthy input sequences. Re-ViLM [55] is tailored for image captioning and it follows the Flamingo architecture and only uses the retrieved text for augmentation. RA-CM3 [56] leverages retrieval to augment the CM3 [2] pipeline which can do both text-to-image and image-to-text generations.
In contrast to previous approaches, our primary objective is to develop a lightweight module capable of processing multimodal context information efficiently and effectively, with superior generalization capabilities. This novel approach enables LLM to proficiently reason from multimodal in-context examples, leading to more accurate and precise inferences. As discussed in [31], LLaVA-1.5 faces limitations when confronted with scenarios involving multiple images and lengthy contexts. CaMML can effectively address these challenges, a fact further corroborated by its superior performances across multiple benchmarks.
3 Context-Aware Multimodal Learner
Figure 2: CaMML framework, which consists of retriever, perceiver and generator. Once receiving user query $q$, CaMML retriever identifies relevant multimodal contexts $C$ from datastore, then CaMML Perceiver seamlessly integrates various modalities, effectively encodeing long-context information and injecting it into the CaMML generator. This allows for the prediction of responses that are conditioned on both the context and the query $q$.
3.1 Architecture of CaMML
The structure of CaMML is depicted in Figure 2. We provide a detailed explanation of each component below.
3.1.1 Datastore and Context Retriever
The datastore is created from either the training set or external resources. We adopt an embedding encoder, ImageBind [15], to extract dense vector representations for every multimodal sample contained in the datastore. Following this, we build an index using Faiss [22], a highly efficient similarity search library, to enable rapid and efficient search operations. During both the training and inference phases, our approach involves identifying the top $N$ most closely related multimodal samples, denoted as $C_{1},C_{2},\ldots,C_{N}$, from the datastore using Faiss, based on a given multimodal query $q$ [^1] Specifically, we compute the similarity between query text and context image, and top-k images along with their corresponding texts are recalled from the datastore.. It is worth noting that the retriever remains frozen throughout all stages of the process.
3.1.2 Multimodal CaMML Perceiver
To seamlessly integrate interleaved multimodal samples, we introduce a novel module denoted as CaMML Perceiver, which revolves around two key design principles: (1) The module should accommodate a dynamic number of context samples without imposing a significant computational burden. It is worth noting that directly concatenating the images tokens and text tokens will result in $L=\sum^{N}_{i=1}(T_{i}^{\text{IMG}}+T_{i}^{\text{TXT}})$ tokens, where $T_{i}^{\text{IMG}}$ is the number of image tokens and $T_{i}^{\text{TXT}}$ is the number of text tokens. The formulation results in a linear increase of $L$ with respect to $N$, posing scalability concerns; (2) It is crucial that the text and image in each individual sample remains tightly coupled.
To achieve this, we embrace a hierarchical design, facilitating the transformation of the $N$ context samples into $M$ tokens, where $M<L$. Formally, the multimodal CaMML Perceiver comprises three key modules: a Vision Perceiver (VP) denoted as $f_{\theta^{\text{VP}}}$ , a Language Perceiver (LP) denoted as $f_{\theta^{\text{LP}}}$, and a Context Perceiver (CP) denoted as $f_{\theta^{\text{CP}}}$. Each of these modules follows a similar architectural pattern with Perceiver Layers as the core component, as depicted in Figure 2.
Vision Perceiver
The Vision Perceiver takes the image feature extracted with ViT, $\mathbf{V}^{(C_{i})}\in\mathbb{R}^{T_{i}^{\text{IMG}}\times d}$, as input, and it undergoes a sequence of transformations. It starts with a self-attention layer to capture inherent visual relationships. Subsequently, cross-attention layers are employed to interact with the text token embeddings $\mathbf{U}^{(C_{i})}$, allowing the model to enrich its understanding by integrating information from the text domain. Finally, the processed data passes through a feed-forward layer, resulting in an output denoted as $f_{\theta^{\text{VP}}}(\mathbf{V}^{(C_{i})})\in\mathbb{R}^{T_{i}^{\text{IMG}}\times d},i=1,...,N$. Subsequently, we concatenate the outputs generated by the Vision Perceiver for each of the $N$ context samples, yielding a feature matrix of shape $\mathbb{R}^{(\sum^{N}_{i=1}T_{i}^{\text{IMG}})\times d}$
$\mathbf{E}^{\text{IMG}}=\{f_{\theta^{\text{VP}}}(\mathbf{V}^{(C_{1})},\mathbf{U}^{(C_{1})}),...,f_{\theta^{\text{VP}}}(\mathbf{V}^{(C_{N})},\mathbf{U}^{(C_{N})})\},$
[…]
Language Perceiver
The Language Perceiver, on the other hand, takes as input the text token embeddings $\mathbf{U}^{(C_{i})}\in\mathbb{R}^{T_{i}^{\text{TXT}}\times d}$, which encapsulate the linguistic content of the sample. The processing pipeline mirrors that of the Vision Perceiver, starting with a self-attention mechanism to capture textual relationships, followed by interactions with the visual embeddings $\mathbf{V}^{(C_{i})}$ via a cross-attention layer to create a holistic understanding of the multimodal context. The final step involves a feed-forward layer, resulting in an output denoted as $f_{\theta^{\text{LP}}}(\mathbf{U}^{C_{i}})\in\mathbb{R}^{T_{i}^{\text{TXT}}\times d}$. Similarly, we concatenate the outputs of all Language Perceivers to obtain a feature matrix of shape $\mathbb{R}^{(\sum^{N}_{i=1}T_{i}^{\text{TXT}})\times d}$:
$\mathbf{E}^{\text{TXT}}=\{f_{\theta^{\text{LP}}}(\mathbf{U}^{(C_{1})},\mathbf{V}^{(C_{1})}),...,f_{\theta^{\text{LP}}}(\mathbf{U}^{(C_{N})},\mathbf{V}^{(C_{N})})\}.$
Context Perceiver
The Context Perceiver take the lengthy outputs from either the Vision Perceiver or the Language Perceiver as input and produces a condensed set of contextual representations. This is achieved by using a fixed number of learnable embeddings of size $\mathbb{R}^{\frac{M}{2}\times d}$ as the input of the Context Perceiver, where $M$ is typically significantly smaller than $L$. These embeddings undergo processing through a self-attention layer, followed by a cross-attention layer with the feature matrix from either the Vision Perceiver or the Language Perceiver, and a feedforward layer. Specifically, for the Vision Perceiver, the learnable embeddings are denoted as $\mathbf{H}^{\text{IMG}}\in\mathbb{R}^{\frac{M}{2}\times d}$, and the output after applying the Context Perceiver is denoted as:
$\tilde{\mathbf{H}}^{\text{IMG}}=f_{\theta^{\text{CP}}}(\mathbf{H}^{\text{IMG}},\mathbf{E}^{\text{IMG}})\in\mathbb{R}^{\frac{M}{2}\times d}.$
Similarily, for the Language Perceiver, the learnable embeddings are denoted as $\mathbf{H}^{\text{TXT}}\in\mathbb{R}^{\frac{M}{2}\times d}$, and the output after applying the Context Perceiver is denoted as:
$\tilde{\mathbf{H}}^{\text{TXT}}=f_{\theta^{\text{CP}}}(\mathbf{H}^{\text{TXT}},\mathbf{E}^{\text{TXT}})\in\mathbb{R}^{\frac{M}{2}\times d}.$
Following this, $\tilde{\mathbf{H}}^{\text{IMG}}$ and $\tilde{\mathbf{H}}^{\text{TXT}}$ are concatenated, resulting in a contextual feature matrix $\tilde{\mathbf{H}}\in\mathbb{R}^{M\times d}$. This matrix is prepended to the beginning of the sequence, along with the text token embeddings of $q$ and the image embeddings of $q$ (transformed by the LLaVA multimodal projector). This integrated data serves as the input for the Large Language Model (LLM) to facilitate further processing.
3.2 Model Training
In the training process, we freeze the retriever and vision encoder, and train CaMML by minimizing the following causal language modeling loss.
$\ell=-\sum_{i=1}^{|y|}\log p_{\theta}(y_{i}|\hat{y}_{1:i-1},q,C_{1},...,C_{N}),$
where $\theta\leftarrow(\theta^{\text{LLM}},\theta^{\text{VP}},\theta^{\text{LP}},\theta^{\text{CP}},\mathbf{H})$ is the model trainable parameter ($\theta^{\text{LLM}}$ is the the parameter of LLM); $y_{i}$ is the ground-truth target, and $\hat{y}_{1:i-1}$ is the $i-1$ preceding tokens of output $y_{i}$; Specifically, we use Vicuna-7B and Vicuna-13B as our backbone language model, resulting in two models CaMML-7B and CaMML-13B.
4 Experiment
Table 1, Table 2, and Table 3 provide a comprehensive summary of the performance achieved by CaMML-7B and CaMML-13B across various multimodal benchmarks. In comparison to previous state-of-the-art approaches, CaMML exhibits noticeable improvement and establishes a new state-of-the-art. We elaborate on the experimental settings in the following sections and offer a comprehensive description (e.g., hyperparamter configuration) of the settings in the supplementary material.
4.1 Multimodal Reasoning on ScienceQA
Method AVG. IMG TXT
Human Average 88.40 87.50 89.60
UnifiedQA [36] 74.11 66.53 66.42
GPT-3 CoT [36] 75.17 67.43 74.68
GPT-4 CoT [38] 83.99 71.49 82.65
LLaMA-Adapter [61] 85.19 80.32 83.72
[…]
LLaVA-7B [32] 89.28 87.32 90.96
LLaVA-13B [32] 90.90 88.00 89.49
CaMML-7B 91.32 89.24 93.21
CaMML-13B 92.03 89.94 93.84
Table 1: Comparison with state-of-the-art methods on ScienceQA benchmark: CaMML finetuned on train split and evaluated on test split. “AVG.” represents the average accuracy of all ScienceQA questions. “IMG” refers to the questions that include image contexts, while “TXT” refers to the questions without any images.
[…]
IDEFICS-9B [25] LLaMA-7B 353M+1M 50.9 38.4 35.5 - 25.9 - - 48.2 25.2 - -
InstructBLIP [12] Vicuna-7B 129M+1.2M - 49.2 34.5 60.5 50.1 - - 36.0 23.7 53.4 26.2
Qwen-VL [6] Qwen-7B 1.4B+50M 78.8 59.3 35.2 67.1 63.8 - - 38.2 7.4 56.3 -
Qwen-VL-Chat [6] Qwen-7B 1.4B+50M 78.2 57.5 38.9 68.2 61.5 - 1487.5 60.6 56.7 58.2 -
LLaVA-1.5 [31] Vicuna-7B 558K+665K 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.3 58.6 30.5
CaMML-7B Vicuna-7B ${}^{\dagger}$+665K 79.4 62.7 51.2 67.9 58.0 86.4 1506.9 66.9 60.6 60.4 32.2
IDEFICS-80B [25] LLaMA-65B 353M+1M 60.0 45.2 36.0 - 30.9 - - 54.5 38.1 - -
BLIP-2 [28] Vicuna-13B 129M 41.0 41.0 19.6 61.0 42.5 85.3 1293.8 - - 46.4 22.4
InstructBLIP [12] Vicuna-13B 129M+1.2M - 49.5 33.4 63.1 50.7 78.9 1212.8 - - - 25.6
Shikra [8] Vicuna-13B 600K+5.5M 77.4 - - - - - - 58.8 - - -
LLaVA-1.5 [31] Vicuna-13B 558K+665K 80.0 63.3 53.6 71.6 61.3 85.9 1531.3 67.7 63.6 61.6 35.4
CaMML-13B Vicuna-13B ${}^{\dagger}$+665K 80.2 63.7 57.4 72.3 59.9 86.7 1588.7 70.2 63.6 62.3 36.4
[…]
ScienceQA, as introduced in [36], serves as a multimodal benchmark designed for the task of science question answering. It comprises 21,000 multiple-choice questions, encompassing various domains including biology, physics, chemistry, Earth science, and more. We use the ScienceQA train split to build the datastore search index and for the model training. In specific, Vicuna-v1.3 is used as the LLM backbone, and CLIP-ViT-L-14 [43] is adopted as the visual feature encoder. By default, we retrieve three samples from the datastore as the contexts. We conduct evaluation of the ScienceQA on test split and adopt the same datastore (i.e., train split) search index for contextual samples retrieval.
[…]
Method (shots) COCO Caption Flickr30k OKVQA VQAv2 Vizwiz
CIDEr CIDEr Acc Acc Acc
Retrieval-augmented models:
 RA-CM3 [56] (2) 89.1 - - - -
 ReViLM [55] (0) 60.8 52.1 - - -
 ReViLM [55] (2) 77.2 - - - -
 ReViLM [55] (4) 90.5 - - - -
 ReViLM [55] (8) 90.2 - - - -
Zero/Few-shot models:
 Uni-Perceiver (0) [67] 109.8 41.2 - - -
 Flamingo-9B (4) 93.1 72.6 49.3 56.3 34.9
 Flamingo-9B (32) 106.3 72.8 51.0 60.4 44.0
 Flamingo-80B (4) 103.2 75.1 57.4 63.1 39.6
 Flamingo-80B (32) 113.8 75.4 57.8 67.6 49.8
 KOSMOS-1 (4) [17] 101.7 75.3 - 51.8 35.3
 MMICL (4) [64] - 72.0 - 70.6 50.3
CaMML-7B (3) 111.4 82.7 64.7 79.4 51.2
CaMML-13B (3) 116.8 84.5 66.3 80.2 57.4
Table 3: Comparison with state-of-the-art zero/few-shot models including retrieval-augmented counterparts on Captioning and VQA tasks: CaMML models are the ones that trained in instruction tuning (Sec 4.2) and CaMML are evaluated with 3-shots. CaMML-13B achieves the best performance on 5/5 tasks, even outperforming Flamingo-80B model [3] under 32 shots.
4.2 Multimodal Instruction Tuning
[…]
We compare CaMML with large multimodal baselines such as BLIP-2 [28], InstructBLIP [12], Shikra [8], IDEFICS-9B [25], IDEFICS-80B [25], Qwen-VL [6], Qwen-VL-Chat [6], and LLaVA-1.5 [31]. As shown in Table 2, CaMML-13B demonstrates the best performance across all 11 benchmarks, surpassing LLaVA-1.5, which uses the same training data. Notably, when employing Vicuna-13B as its backbone, CaMML-13B outperforms BLIP-2, InstructBLIP-13B, Shikra, and LLaVA-1.5-13B. Additionally, CaMML-7B delivers impressive results, outperforming all other baseline models except for LLaVA-1.5-13B on ten of the benchmark datasets. On POPE, CaMML-7B achieves performance on par with that of CaMML-13B. These encouraging observations showcase CaMML’s efficacy in harnessing multimodal contextual information and following multimodal instructions. In contrast to Liu et al. [31]’s assertion that LLaVA-1.5 cannot handle multiple images due to the absence of instruction-tuning data and context length limitations, our design demonstrates feasibility of training a model to effectively process multiple images merely with the same training set and efficiently manage long context length.
4.3 Multimodal Few-shot Learning
CaMML essentially operates as a few-shot learning method. Few-shot models leverage the synergy of multiple intertwined images and texts to guarantee precise inference. Among these, retrieval-augmented models are widely recognized for their ability to retrieve highly similar samples, thereby improving few-shot prediction. In Table 3, we conduct a comparison between CaMML, retrieval-augmented models (RA-CM3 [56] and ReViLM [55]), and other popular few-shot LLMs such as Flamingo [3], KOSMOS-1 [17], MMICL [64] on tasks including visual captioning (COCO caption, Flickr30k), and visual question answering (OKVQA, VQAv2, Vizwiz). For CaMML, we employ the same model described in Section 4.2, with LLaVA-665K serving as the datastore.
Several observations merit attention: (1) It is worth mentioning that while previous visual language models augmented with retrieval capabilities have been predominantly optimized for visual captioning tasks, CaMML exhibits broad applicability across a wide range of tasks, including visual question answering and visual captioning. (2) Notably, CaMML-13B outshines the significantly larger Flamingo-80B, highlighting CaMML’s superior efficiency and architectural effectiveness despite its smaller size. (3) Remarkably, CaMML, even with a mere three-shot setup, surpasses models that necessitate 32 shots for equivalent tasks, underscoring its exceptional data efficiency and learning prowess.
Method AVG. IMG TXT
CaMML-7B Baseline 91.3 89.2 93.2
 - w/o Perceiver 89.7 85.8 93.2
 - w/o Vision Perceiver 89.8 87.4 92.1
 - w/o Language Perceiver 90.0 87.7 92.0
 - w/o Shared weights 91.3 89.1 93.2
Table 4: Ablation Experiments on CaMML perceiver components and the existence of shared-weights Context Perceiver: CaMML-7B alternatives are evaluated on ScienceQA-test.
5 Model Analyses
5.1 Quantitative Analyses
Here, we conducted an analysis of CaMML, using the CaMML-7B model for expeditious experiments, on the ScienceQA dataset. Our goal was to discern the significance of each module within the CaMML architecture, as well as the influence of critical hyperparameters such as the number of layers, the selection of $M$, and the quantity of retrieved context samples. The default settings for CaMML-7B baseline are as follows: the number of layers is set to 2 for all perceivers, $M=128$, $N=3$, and hidden size is set to 768.
5.1.1 Contribution of each components
We conducted ablation studies on the Perceiver, Vision Perceiver, Language Perceiver and Context Perceiver (shared weights or not) within CaMML to evaluate the individual contributions of these components. The outcomes are documented in Table 4. Our results highlight the critical role played by each component. The removal of Perceivers, in particular, leads to a marked deterioration in performance, indicating its significant influence on the model’s overall effectiveness. Also, we have observed nearly identical results for both alternatives of the Context Perceiver, whether the weights are shared or not. Therefore, we prefer selecting the shared-weights option to save on computation.
Figure 3: Ablation Experiments on CaMML perceiver hyper-parameters: layers, query number $M$ and hidden sizes. CaMML-7B with different settings are evaluated on ScienceQA test.
Figure 4: Ablation Experiments on CaMML context number $N$. Left: different CaMML models trained on $N$ shots are evaluated under 1$\sim$32 shots. Right: comparison between CaMML and CaMML without perceiver in terms of inference running time and memory footprint, the statistic is averaged on 100 samples from CaMML-7B, which are tested on NVIDIA A100-80G GPU using ScienceQA dataset.
5.1.2 Impact of Hyperparameters
We conducted a thorough exploration of the influence of hyperparameters on the performance of CaMML.
In Figure  3, we report the performance by varying the number of layers, hidden sizes, and query number $M$. Our observations reveal several key insights: (1) Increasing the number of layers in CaMML can have a detrimental effect on model performance. While we observe similar performance with both 2 layers and 12 layers, we opt for 2 layers due to its smaller model size; (2) Increasing the value of $M$ does not consistently lead to performance improvements; (3) Larger hidden sizes tend to yield more favorable results in our analysis.
Furthermore, we investigated the impact of the number of context samples, $N$, used in the training stage as well as the inference stage. We conduct experiments where we varied the value of $N$ during training and inference, and the corresponding performance is reported in Figure 4. We observe that: (1) It is easy for CaMML to accommodate a large number of shots; (2) Increasing the value of $N$ during inference does not consistently result in improved performance. Similar trends have been reported in previous works such as [3] and [55]. One plausible explanation for this phenomenon could be that a longer context might introduce complexity and potentially convolute the inference process; (3) The “mixed-shots” training strategy (Sec 4.2) has shown the potential to yield constant superior performance when compared to using a fixed value for $N$.
Figure 5: Visualization of context-aware CaMML vs. no-context LLaVA-1.5. Left: sketch drawing of the Great Wall. Right: depiction of the biology metamorphosis of a butterfly.
Figure 6: Visualization of CaMML vs. GPT-4V &\&& LLaVA-1.5. CaMML demonstrates a strong understanding of contexual sequences. We directly input contexual sequences (Left: consecutive Homer Simpson video frames and Right: Lionel Messi winning Ballon d’or in October 2023.) to CaMML perceiver without additional contexts.
5.1.3 Inference Cost
As outlined in the model design, our context model adeptly manages large samples by condensing a significant number of raw tokens into a more streamlined representation. This efficient process leads to a faster forward pass in subsequent LLMs. We compared CaMML with a baseline approach, where all context tokens are directly input into LLMs without CaMML perceivers, and we present the inference speed and memory footprint in Figure 4 (right). We have noticed that CaMML only incurs negligible additional cost when increasing the number of context sample; As the number of context samples grows, the efficiency and memory benefits of CaMML become more significant compared with the baseline method.
5.2 Qualitative Analysis
5.2.1 The Importance of Context-Awareness
CaMML effectively processes a wide range of contextual inputs. In Figure 5, we compare the response from CaMML with context sample support and LLaVA-1.5, emphasizing the importance of context-awareness. The comparison illustrates that the inclusion of relevant context examples is crucial for ensuring the accuracy of answers. CaMML can not only infer using highly relevant context samples (The Great Wall example), but also draw insights from analogous domain-specific samples (The Metamorphosis example) and conduct multimodal analogy-based learning [5].
5.2.2 Handling Image Sequences
We also demonstrate that CaMML possesses the ability to handle image sequences, even in the absence of explicit training for this specific task. In Figure 6, instead of retrieval, we consider the last image of the image sequence as the $q$ and the preceding images as context samples. A comparison with LLaVA-1.5 and GPT-4V in Figure 6 on the QA task on image sequences reveals that while LLaVA-1.5 and GPT-4V struggles to fully comprehend the image sequences and learn from the priors, resulting in incorrect answers, CaMML can effectively understand the image sequences and give accurate responses. More importantly, CaMML can provide accurate inferences given access to the most up-to-date information regardless of LLM’s internal knowledge, yet both LLaVA-1.5 and GPT-4V lack the capability to incorporate real-time updates of world knowledge.
Figure 7: Visualization of CaMML vs. GPT-4V &\&& LLaVA-1.5. CaMML effectively mitigates hallucinations and accurately identifies objects. On the left, there is an illusion of muffin and chihuahua. On the right, there is an illusion of fried chicken and labradoodle.
5.2.3 Tackling Multimodal Hallucination
Previous research indicates that RAG is effective in reducing hallucinations in content generation [48]. Here, we aim to investigate if CaMML can handle hallucinations in a multimodal setting. Figure 7 illustrates a multimodal QA task involving the recognition of similar objects: Muffins and Chihuahuas; Labradoodles and Fried Chicken. We treat the grid of images as a single image. It is evident that LLaVA-1.5 and GPT-4V face challenges in distinguishing between these objects, leading to hallucinated answers. In contrast, CaMML accurately answers the questions with the groundness of relevant context samples.
6 Conclusion
We present a new methodology called CaMML, a context-aware multimodal learner designed for the fine-tuning of large multimodal models. With CaMML, we build two multimodal models, CaMML-7B and CaMML-13B. CaMML empowers them to draw insights from analogous, domain-specific, and up-to-date context samples to make grounded inferences. Moreover, it employs a lightweight multimodal perceivers to seamlessly integrate these context samples, enabling an efficient processing of lengthy context tokens. Our proposed CaMML-13B model achieves state-of-the-art results across more than ten prominent multimodal benchmarks, surpassing previous methods by a substantial margin. These achievements underscore CaMML’s effectiveness in various multimodal applications.
Appendix A Extended Experiments
A.1 Experimental Setup
Summary of Datasets and Benchmarks.
ScienceQA [36] train split is used for ScienceQA finetuning and ablation study. In our instruction-tuning training, we adopt LLaVA-665K dataset, which contains LLaVA-158K, ShareGPT-40k [1], VQAv2 [4], GQA [18], OKVQA [39], OCRVQA [40], A-OKVQA [46], TextCaps [49], RefCOCO [58] and VG [24]. In our evaluation, VQAv2, GQA, TextVQA [50], MME [14], POPE [29], MM-Vet [59], ScienceQA, MMBench [34], MMBench-CN [34], SEED-Bench [27] and Vizwiz [16] are considered as benchmarks. We also conduct evaluation on COCO Caption [10], Flickr30k [41], OKVQA [39], A-OKVQA [46], and RefCOCO/+/g [58]. In our qualitative visualization, we adopt another 2M datastore as the source of context examples, which comprised from external resources incorporates 2,348K multimodal samples, ranging from captioning with BLIP-LAION’s 558K entries and Local Narrative [42], to knowledge-based QA with KVQA [45], narrative-driven QA from VCR [60] and Visual7W [68], visual grounding via RefCOCOPlus [58] and RefCOCOg [58], OCR from TextOCR [51], along with the in-domain LLaVA-665K set.
Summary of Pretrained Checkpoints.
We utilize Vicuna-7B/13B [65] as our foundation Large Language Model (LLM), ViT-L-14 architecture is used as the vision encoder. In detail, Vicuna-v1.3333https://huggingface.co/lmsys/vicuna-7b-v1.3, https://huggingface.co/lmsys/vicuna-13b-v1.3 and CLIP-ViT-L-14444https://huggingface.co/openai/clip-vit-large-patch14 [43] initialization is for ScienceQA finetuning and Ablation studies, Vicuna-v1.5555https://huggingface.co/lmsys/vicuna-7b-v1.5, https://huggingface.co/lmsys/vicuna-13b-v1.5, CLIP-ViT-L-14-336px [^6] https://huggingface.co/openai/clip-vit-large-patch14-336, and LLaVA-1.5 multimodal projector [^7] https://huggingface.co/liuhaotian/llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, https://huggingface.co/liuhaotian/llava-v1.5-mlp2x-336px-pretrain-vicuna-13b-v1.5 is initialized for instruction finetuning. For the source of contexts, the ImageBind [^8] https://dl.fbaipublicfiles.com/imagebind/imagebind_huge.pth [15]-Huge model is adopted to embed texts and images, computing their similarities for indexing top-k samples.
Implementation Details.
 [^9] All the experiments are trained under Deepspeed Zero-3 FP16 configuration.
In our ScienceQA finetuning, we train for 12 epochs, with batch size 4 per GPU and learning rate 2e-5. We illustrate the hyperparameter configurations here for ScienceQA finetuning in main paper Table 1:
  • CaMML-7B: {number of query $M$=128, number of perceiver layer 2, perceiver layer hidden size 768, number of shots $N$=1},
  • CaMML-13B: {number of query $M$=256, number of perceiver layer 2, perceiver layer hidden size 4608, number of shots $N$=3}.
In our instruction finetuning, we train for 1 epoch on 8GPUs, with batch size 8 per GPU and learning rate 2e-5, the LLM is activated as well as CaMML perceivers, while the vision encoder and CaMML retriever is frozen. We illustrate the hyperparameter configurations here for instruction finetuning in main paper Table 2&3:
  • CaMML-7/13B: {number of query $M$=128, number of perceiver layer 2, perceiver layer hidden size 768, mixed-training shots 1$\sim$3 and inferenced shots $N$=3}.
A.2 CaMML Retrieval Setup
We built CaMML retriever in following steps:
  • Utilize ImageBind [15] model to inference upon images and get corresponding visual embedding.
  • Utilize Faiss [21] to build datastore index for visual embedding, and bind each data source (image and text) to the index, ensuring the right one-to-one retrieval.
  • For each query, we forward with ImageBind model and compute similarity between query embedding and datastore index. We select top-k similarity samples as our contexts.
Note that, in our quantitative experiments, we compute query (text) embedding’s similarity with (vision) datastore, while in qualitative analyses, we compute query (vision) embedding’s similarity with (vision) datastore, to obtain relevant image information.
A.3 Experimental Results
Method (shots) A-OKVQA RefCOCO RefCOCO+ RefCOCOg
Acc Acc Acc Acc
CaMML-7B (3) 81.1 66.6 60.3 57.6
CaMML-13B (3) 82.0 70.6 65.9 60.5
Table 5: CaMML multimodal performance on A-OKVQA, Refcoco/+/g.
A.3.1 Multimodal Task Performance
In addition to the results presented in the main paper tables, CaMML demonstrates versatility in handling various multimodal tasks without requiring further fine-tuning. Table 5 shows that CaMML achieves exceptional performance on the Augmented Outside-Knowledge VQA (A-OKVQA) task with an accuracy of 82.0, and also exhibits good ability in localizing visual grounding tasks such as RefCOCO/+/g.
A.3.2 Finegrained Evaluation
CaMML is tested on MMBench and MMBench-CN to showcase each fine-grained ability such as Logic Reasoning (LR), Attribute Recognition (AR), RR (Relation Reasoning), Instance-Level Fine-Grained Perception (FP-S), Cross-Instance Fine-Grained Perception (FP-C), Coarse Perception (CP). According to Table  6, CaMML-13B has achieved 8 out of 12 state-of-the-art results among large multimodal models.
Method LLM MMBench-dev MMBenchCN-dev
Overall LR AR RR FP-S FP-C CP Overall LR AR RR FP-S FP-C CP
MMICL [64] FLANT5-XXL 67.9 49.2 71.6 73.0 66.7 57.2 77.2 - - - - - - -
mPLUG-Owl2 [57] LLaMA2-7B 66.5 32.2 72.4 60.9 68.6 60.1 79.4 59.5 28.8 64.8 48.7 60.1 50.3 76.0
SPHINX [30] LLaMA2-13B 67.2 33.1 67.3 58.3 74.4 59.4 80.7 58.6 21.2 61.8 43.5 62.1 58.7 73.6
LLaVA-1.5 [31]-7B Vicuna-7B 63.0 26.3 68.8 53.0 67.2 56.6 76.4 57.4 25.4 58.8 55.7 55.3 49.7 75.7
LLaVA-1.5 [31]-13B Vicuna-13B 68.2 44.1 67.3 60.0 72.0 59.4 82.1 61.9 36.4 65.8 49.6 62.1 59.4 75.0
CaMML-7B Vicuna-7B 66.9 34.2 71.6 66.1 70.4 56.5 78.9 60.6 27.1 87.0 55.7 58.0 57.3 77.0
CaMML-13B Vicuna-13B 70.2 44.2 87.5 60.9 72.4 67.8 84.6 63.5 37.3 88.9 52.2 63.1 59.4 77.0
Table 6: Experiments on fine-grained multimodal reasoning ability: CaMML evaluated on MMBench and MMBenchCN, compared with other state-of-the-art methods. The categories include Logic Reasoning (LR), Attribute Recognition (AR), RR (Relation Reasoning), Instance-Level Fine-Grained Perception (FP-S), Cross-Instance Fine-Grained Perception (FP-C), Coarse Perception (CP).
Appendix B Additional Cases
[…]
Figure 10: CaMML capability of localization on implicit description.
Image Generation
CaMML demonstrates its ability to generate reliable content based on a wealth of multimodal contextual sources. In Figure  8 9, we present an example of CaMML’s capability in prompting image generation. In the case, CaMML is known for its ability to incorporate contextual elements into generating descriptions. For example, it can generate descriptions of traditional Chinese-style buildings in cyber-punk style. When using CaMML, the generated images are likely to include a mix of old buildings with a modern style, based on the provided contextual samples. On the other hand, when using simple prompt without any contextual samples, the generated image is more likely to be limited to modern style skyscrapers only.
Implicit-Description Localization
CaMML also highlights its strong localization capabilities, particularly in terms of reasoning the concept behind and high-level understanding. We utilize CaMML to detect objects by providing implicit descriptions and allowing CaMML to reason about the nature and location of these objects. As shown in Figure 10, CaMML is able to identify the objects with implicit descriptions (such as airplane: “The thing that carries hundreds of souls to the sky” or mirror: “The thing that makes two same dogs existing in this image”, etc..).